OpenStudy (anonymous):

For every two feet the ramp goes up it also goes 5 feet forward. If this ratio remained constant, how many feet high would a ramp be if it was 22.5 feet long?

OpenStudy (anonymous):

Ratio: 2 feet h/5 feet long Therefore: 2/5 = x / 22.5 x = 9